If the distance travelled by a body in nth second is (4+6n)m then find the initial velocity and acceleration of the body respectively.

Grade:9

1 Answers

CHALLA NAVNEESWAR REDDY
7 years ago
given s=(4+6n) at t=n sec
we know that, v=ut+1/2at2
=(4+6n)=un+11/2an2,  u=4+6n-1/2n2,at t=0, initial velocity(u) =4m/sec
a=(4+6n-un)2/t2, at t=0, acceleration(a)=0m/sec2